OpenStudy (anonymous):

solve cos^2 x -3sinx=3

OpenStudy (anonymous):

start with \[1-\sin^2(x)-3\sin(x)-3=0\] and solve the quadratic equation in \(\sin(x)\)

OpenStudy (jhannybean):

Oh nevermind, you mentioned that already... Then you can replace all the sine's with y's

OpenStudy (jhannybean):

\[1-y^2-3y-3=0\] reduces to \[-y^2-3y-2=0\]